# DRAL: Deep Reinforcement Adaptive Learning for Multi-UAVs Navigation in Unknown Indoor Environment

### v1

@@ Abstract
Autonomous indoor navigation of UAVs presents numerous challenges, primarily due to the limited precision of GPS in enclosed environments. Additionally, UAVs’ limited capacity to carry heavy or power-intensive sensors, such as overheight packages, exacerbates the difficulty of achieving autonomous navigation indoors. This paper introduces an advanced system in which a drone autonomously navigates indoor spaces to locate a specific target, such as an unknown Amazon package, using only a single camera. Employing a deep learning approach, a deep reinforcement adaptive learning algorithm is trained to develop a control strategy that emulates the decision-making process of an expert pilot. We demonstrate the efficacy of our system through real-time simulations conducted in various indoor settings. We apply multiple visualization techniques to gain deeper insights into our trained network. Furthermore, we extend our approach to include an adaptive control algorithm for coordinating multiple drones to lift an object in an indoor environment collaboratively. Integrating our DRAL algorithm enables multiple UAVs to learn optimal control strategies that adapt to dynamic conditions and uncertainties. This innovation enhances the robustness and flexibility of indoor navigation and opens new possibilities for complex multi-drone operations in confined spaces. The proposed framework highlights significant advancements in adaptive control and deep reinforcement learning, offering robust solutions for complex multi-agent systems in real-world applications.
@@ I Introduction
Micro Aerial Vehicles (MAVs), such as quadcopters outfitted with cameras, have become integral to a wide range of applications [1, 2, 3, 4, 5, 6], including search and rescue operations, exploratory missions, and entertainment. While outdoor autonomous navigation has achieved significant success leveraging the global positioning system (GPS), the precision of GPS diminishes substantially in indoor environments, posing substantial challenges for autonomous indoor flight.
@@
Various methodologies have been proposed to address the complexities of indoor autonomous navigation. One prominent approach is Simultaneous Localization and Mapping (SLAM), which employs laser range finders, RGB-D sensors [7], or single cameras to generate a three-dimensional map of unknown indoor environments and ascertain the MAV’s position within this map. Another notable method [8] involves stereo vision, wherein depth perception is achieved by calculating the disparity between stereo images. However, SLAM is often impractical for MAVs due to the high computational demand of constructing a 3D model [9], and the resultant 3D structures tend to underperform in feature-sparse environments such as plain walls. Similarly, depth estimation via stereo vision suffers from poor performance in texture-less regions and issues with specular reflections. Additionally, the prevalence of commercially available quadcopters equipped with only a single camera renders these solutions impractical for widespread use.
Modern environments are primarily designed for human use, making humanoid robots with their human-like skeletal structures particularly well-suited for tasks in these settings. [10, 11, 12, 13], a novel method with the SVM model has been proposed to pave the path to solving the issues in this field, which shows good performance in complicated and changing scenarios. Moreover, [14] developed a novel algorithm for a Monte Carlo tree search algorithm, which improved the searching efficiency for robotics decision tasks. These robots offer distinct advantages over other types of robots in human-centric environments. Recently, massively parallel deep reinforcement learning [15, 16, 17] in simulation has gained popularity. However, due to the intricate structure of humanoid robots, the sim-to-real gap is more pronounced compared to that of quadrupedal robots. One critical challenge in achieving synchronization within drone networks stems from the inevitable variations in their dynamics and parameters [18, 19]. These discrepancies arise from manufacturing inconsistencies, wear and tear, environmental factors, and differences in payloads, presenting substantial obstacles to developing effective control strategies for the network as [20]. The disparity between the behaviors exhibited by agents in simulation versus those observed in real-world scenarios is referred to as the reality gap.
@@
This paper explores the integration of state-of-the-art (SOTA) adaptive control with deep reinforcement learning methods [21, 22, 23, 24] for multi-drone systems, specifically targeting navigation tasks within indoor environments. The contributions of the paper are:
@@ II Related work
Robotic exploration, which leverages multi-robot systems to navigate and map uncharted environments, has been the subject of extensive research. Some studies prioritize rapid spatial exploration [25], aligning with the objectives of this paper, while [26] emphasize precise environmental reconstruction. Among the myriad methodologies proposed, frontier-based approaches stand out as classic techniques. These methods were initially introduced in related works and subsequently subjected to a more thorough evaluation in [27]. A stochastic differential equation-based method was proposed in [28] to identify frontiers in three-dimensional space. The closest frontier was chosen as the next target in the seminal work. Conversely, [29] presented an alternative strategy that, during each decision-making process, selects the frontier within the field of view that minimizes velocity changes, thereby maintaining a consistently high flight speed. This strategy has been demonstrated to outperform the traditional method [30, 31, 32, 33]. Furthermore, in [34], a differentiable measure of information gain based on frontiers was introduced, enabling path optimization through gradient information.
@@ IV Experiment Results
In our experimental setup, we employ three drones to elevate an unidentified box within a simulation environment modeled on an indoor room constructed in Gazebo, as illustrated in Fig. 2. The local grid maps encompass an area of $14m\times 8m$. Each drone is stationed at the center of its respective local map and is confined to perceiving information within a 3-meter radius, consequently constraining its long-range planning capabilities. These parameters are uniformly maintained across both simulated and real-world experimental conditions.
We evaluate our proposed framework through simulation in Gazabo, benchmarking it in Q-learning, Deep Q learning Sarsa and our method to compare each other. In all tests, the dynamic limits are uniformly set to $v_{\text{max}}=3.4$ m/s and $\omega_{\text{max}}=0.57$ rad/s for all methods. The sensors’ fields of view are configured to $[50\times 50]\ \text{deg}$ with a maximum range of 3 meters. Each method is executed three times in both scenarios, maintaining the same initial configuration. The four methods’ statistical outcomes and exploration progress are presented in Tab. I and 4, respectively. Tab. I shows that our method shows a higher success rate than the other three popular methods in the different transportation tasks during indoor navigation, which means the success rate for three drones carrying an unknown payload can transport from the initial position to the target position without any collision and crush is higher than other methods’ verifications.

### v2

@@ Abstract
Autonomous indoor navigation of UAVs presents numerous challenges, primarily due to the limited precision of GPS in enclosed environments. Additionally, UAVs’ limited capacity to carry heavy or power-intensive sensors, such as overheight packages, exacerbates the difficulty of achieving autonomous navigation indoors. This paper introduces an advanced system in which a drone autonomously navigates indoor spaces to locate a specific target, such as an unknown Amazon package, using only a single camera. Employing a deep learning approach, a deep reinforcement adaptive learning algorithm is trained to develop a control strategy that emulates the decision-making process of an expert pilot. We demonstrate the efficacy of our system through real-time simulations conducted in various indoor settings. We apply multiple visualization techniques to gain deeper insights into our trained network. Furthermore, we extend our approach to include an adaptive control algorithm for coordinating multiple drones to lift an object in an indoor environment collaboratively. Integrating our DRAL algorithm enables multiple UAVs to learn optimal control strategies that adapt to dynamic conditions and uncertainties.
@@ I Introduction
Micro Aerial Vehicles (MAVs), such as quadcopters outfitted with cameras, have become integral to a wide range of applications [1, 2, 3, 4, 5, 6, 7], including search and rescue operations, exploratory missions, and entertainment. While outdoor autonomous navigation has achieved significant success leveraging the global positioning system (GPS), the precision of GPS diminishes substantially in indoor environments, posing substantial challenges for autonomous indoor flight.
@@
Various methodologies have been proposed to address the complexities of indoor autonomous navigation. One prominent approach is Simultaneous Localization and Mapping (SLAM), which employs laser range finders, RGB-D sensors [8], or single cameras to generate a three-dimensional map of unknown indoor environments and ascertain the MAV’s position within this map. Another notable method [9] involves stereo vision, wherein depth perception is achieved by calculating the disparity between stereo images. However, SLAM is often impractical for MAVs due to the high computational demand of constructing a 3D model [10], and the resultant 3D structures tend to underperform in feature-sparse environments such as plain walls. Similarly, depth estimation via stereo vision suffers from poor performance in texture-less regions and issues with specular reflections. Additionally, the prevalence of commercially available quadcopters equipped with only a single camera renders these solutions impractical for widespread applications [11, 12].
Modern environments are primarily designed for human use, making humanoid robots with their human-like skeletal structures particularly well-suited for tasks in these settings. [13], a novel method with the SVM model has been proposed to pave the path to solving the issues in this field, which shows good performance in complicated and changing scenarios. This method offer distinct advantages over other types of robots in human-centric environments. Recently, massively parallel deep reinforcement learning [14, 15, 16] in simulation has gained popularity. However, due to the intricate structure of humanoid robots, the sim-to-real gap is more pronounced compared to that of quadrupedal robots. One critical challenge in achieving synchronization within drone networks stems from the inevitable variations in their dynamics and parameters [17]. For this problem,  [18] presents a novel decentralized adaptive control method that enables a team of drones to cooperatively transport unknown payloads, effectively adapting to uncertainties in mass, inertia, and grasping points in both gravitational and zero-gravity environments, demonstrating robustness even in the event of vehicle loss during the mission. Also, an extension work in [19] makes a significant contribution to aerospace robotics by introducing, for the first time, a novel adaptive detumbling algorithm specifically designed for a non-rigid satellite with unknown dynamic properties, offering a pioneering and high-efficiency solution to address the complexities of flexible satellite stabilization in space, a challenge that had not been fully tackled in previous research. These discrepancies arise from manufacturing inconsistencies, wear and tear, environmental factors [20, 21, 22], and differences in payloads, presenting substantial obstacles to developing effective control strategies for the network as [23]. The disparity between the behaviors exhibited by agents in simulation versus those observed in real-world scenarios is referred to as the reality gap.
@@
This paper explores the integration of state-of-the-art (SOTA) adaptive control with deep reinforcement learning methods [24, 25, 26, 27, 28] for multi-drone systems, specifically targeting navigation tasks within indoor environments. The contributions of the paper are:
@@ II Related work
Robotic exploration, which leverages multi-robot systems to navigate and map uncharted environments, has been the subject of extensive research. Some studies prioritize rapid spatial exploration [29], aligning with the objectives of this paper, while [30] emphasize precise environmental reconstruction. Among the myriad methodologies proposed, frontier-based approaches stand out as classic techniques. Moreover, [31] developed a novel algorithm for a Monte Carlo tree search algorithm, which improved the searching efficiency for robotics decision tasks. These methods were initially introduced in related works and subsequently subjected to a more thorough evaluation in [32]. The closest frontier was chosen as the next target in the seminal work. Conversely, [33] presented an alternative strategy that, during each decision-making process, selects the frontier within the field of view that minimizes velocity changes, thereby maintaining a consistently high flight speed. This strategy has been demonstrated to outperform the traditional method [34, 35, 36, 37]. Furthermore, in [38], a differentiable measure of information gain based on frontiers was introduced, enabling path optimization through gradient information.
@@ IV Experiment Results
In our experimental setup, we employ three drones to elevate an unidentified box within a simulation environment modeled on an indoor room constructed in Gazebo, as illustrated in Fig. 2. The local grid maps encompass an area of $14m\times 8m$. Each drone is stationed at the center of its respective local map and is confined to perceiving information within a 3-meter radius, consequently constraining its long-range planning capabilities as shown in Fig. 3. These parameters are uniformly maintained across both simulated and real-world experimental conditions.
We evaluate our proposed framework through simulation in Gazabo, benchmarking it in Q-learning, Deep Q learning Sarsa and our method to compare each other. In all tests, the dynamic limits are uniformly set to $v_{\text{max}}=3.4$ m/s and $\omega_{\text{max}}=0.57$ rad/s for all methods. The sensors’ fields of view are configured to $[50\times 50]\ \text{deg}$ with a maximum range of 3 meters. Each method is executed three times in both scenarios, maintaining the same initial configuration. The four methods’ statistical outcomes and exploration progress are presented in Tab. I and Fig. 4, respectively. Tab. I shows that our method shows a higher success rate than the other three popular methods in the different transportation tasks during indoor navigation, which means the success rate for three drones carrying an unknown payload can transport from the initial position to the target position without any collision and crush is higher than other methods’ verifications.